### v2

# Colorful positive bases decomposition and Helly-type results for cones

Abstract.
We prove the following colorful Helly-type result: Fix $k\in[d-1]$. Assume $\mathcal{A}_{1},\dots,\mathcal{A}_{d+(d-k)+1}$ are finite sets (colors) of nonzero vectors in ${\mathbb{R}}^{d}$. If for every rainbow sub-selection $R$ from these sets of size at most $\max\{d+1,2(d-k+1)\}$, the system $\left\langle a,x\right\rangle\leq 0,\;a\in R$ has at least $k$ linearly independent solutions, then at least one of the systems $\left\langle a,x\right\rangle\leq 0,\;a\in\mathcal{A}_{i},$ $i\in[d+(d-k)+1]$ has at least $k$ linearly independent solutions.
A rainbow sub-selection from several sets refers to choosing at most one element from each set (color).
The Helly number $\max\{d+1,2(d-k+1)\}$ and the number of colors d+(dk)+1d+(d-k)+1 are optimal.
Our key observation is a certain colorful Carathéodory-type result for positive bases.
Key words and phrases:
Helly-type result, Colorful Helly theorem, Carathéodory lemma, positive basis
2020 Mathematics Subject Classification:
52A35 (primary), 52A30
The author is supported by Projeto Paz and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – 23038.015548/2016-06
1. Introduction
The celebrated Helly’s theorem [Hel23] posits that within a finite family of convex sets in ${\mathbb{R}}^{d}$, if the intersection of any subfamily, consisting of at most $d+1$ sets, shares a common point, then all the sets in the family share at least one common point. This gem of combinatorial convexity has seen many different extensions and generalizations (see [BK22] for one of the most up-to-date surveys). A now-classical result by Lovász (see [Bár82]) is the following “colorful” version of Helly’s theorem:
Proposition 1.1 (Colorful Helly theorem).
Let $\mathcal{F}_{1},\dots,\mathcal{F}_{d+1}$ be finite families of convex sets in ${\mathbb{R}}^{d}.$ Suppose that for any choice $F_{1}\in\mathcal{F}_{1},\dots,F_{d+1}\in\mathcal{F}_{d+1},$ the intersection $\bigcap\limits_{i=1}^{d+1}F_{i}$ is non-empty. Then for some $i\in\left\{1,\dots,d+1\right\},$ the intersection of all sets in the family $\mathcal{F}_{i}$ is non-empty.
Clearly, Lovász’s result implies the original Helly’s theorem by taking $\mathcal{F}_{1}=\dots=\mathcal{F}_{d+1}=\mathcal{F}$.
Recently, Bárány [Bár24, Corollary 4.1] obtained the following Helly-type result, generalizing the result of Katchalski [Kat78]:
Proposition 1.2.
Assume $A$ is a finite set of nonzero vectors in ${\mathbb{R}}^{d}.$ The homogeneous system of linear inequalities $\left\langle a,x\right\rangle\leq 0,$ $a\in A$ has at least $k$ linearly independent solutions if and only if for every $B\subset A$ whose size is at most $\max\{d+1,2(d-k+1)\},$ the system $\left\langle b,x\right\rangle\leq 0,$ $b\in B$ has at least $k$ linearly independent solutions.
Bárány also showed that the constant $m(k,d)=\max\{d+1,2(d-k+1)\}$ is optimal in the following sense. There are examples of finite subsets of ${\mathbb{R}}^{d}\setminus\left\{0\right\}$ such that the solution set of any subsystem corresponding to a subset of at most m(k,d)1m(k,d)-1 points contains $k$ linearly independent solutions, but the solution set of the original system does not.
Suppose one has a finite collection of sets (colors). A rainbow sub-selection from these sets refers to choosing at most one element from each set (color). We note that for convenience, we do not require a rainbow sub-selection to intersect each of the original sets. For the sake of convenience, we will use $[n]$ to denote the set $\{1,\dots,n\}$ for a natural $n.$
Our foremost result is a “colorful” extension of Bárány’s result:
Theorem 1.3.
Fix $k\in[d-1]$. Assume $\mathcal{A}_{1},\dots,\mathcal{A}_{d+(d-k)+1}$ are finite sets of nonzero vectors in ${\mathbb{R}}^{d}.$ If for every rainbow sub-selection $R$ from these sets of size at most $\max\{d+1,2(d-k+1)\}$, the system $\left\langle a,x\right\rangle\leq 0,\;a\in R$ has at least $k$ linearly independent solutions, then at least one of the systems $\left\langle a,x\right\rangle\leq 0,\;a\in\mathcal{A}_{i},$ $i\in[d+(d-k)+1]$ has at least $k$ linearly independent solutions.
Clearly, this theorem implies Proposition 1.2, and the constant $m(k,d)$ is optimal since it is optimal for its “monochromatic” version - Proposition 1.2. Surprisingly, the number of color classes d+(dk)+1d+(d-k)+1 is optimal as well. We find this especially intriguing since usually the Helly numbers for “colorful” and “monochromatic” Helly-type results are the same. Note that for some quantitative colorful Helly-type results [DFN21], the precise Helly number remains unknown.
Notably, when $d=k$, both Proposition 1.2 and Theorem 1.3 can be derived from the Colorful Helly theorem (Proposition 1.1) with $m(d,d)=d+1.$ This observation is proposed as an exercise for the reader.
Our proof of Theorem 1.3 mostly follows Bárány’s proof of Proposition 1.2 with a critical distinction lying in the application of what we term the “Colorful Reay’s decomposition.”
Let us recall another gem of combinatorial convexity, namely the “Colorful Carathéodory theorem” obtained by Bárány in [Bár82]:
Proposition 1.4 (Colorful Carathéodory theorem).
Assume $S_{1},\dots,S_{d+1}$ are subsets of ${\mathbb{R}}^{d}.$ If a point $p$ belongs to the convex hull of each of the sets $S_{1},\dots,S_{d+1},$ then there are points $s_{1}\in S_{1},\dots,s_{d+1}\in S_{d+1}$ such that $p$ belongs to the convex hull of $\left\{s_{1},\dots,s_{d+1}\right\}.$
We introduce a key observation regarding positive bases: a “colorful” extension of Reay’s theorem [Rea65, Theorem 2.6]. This can be seen as a Colorful Carathéodory theorem for positive bases.
First, we recall some notions. Let $S$ be a subset of a linear space. Its positive hull, denoted $\mathrm{pos}\ \!S,$ is the set of all finite linear combinations of elements of $S$ with non-negative coefficients. In other words, $\mathrm{pos}\ \!S=\bigcup\limits_{\lambda\geq 0}\lambda\,\mathrm{conv}S,$ where $\mathrm{conv}S$ stands for the convex hull of $S.$ A set $B$ forms a positive basis of a finite-dimensional linear space $L$ if $L=\mathrm{pos}\ \!B$ and removing any element from $B$ changes the positive hull, meaning $L=\mathrm{pos}\ \!B$, yet $L$ differs from $\mathrm{pos}\ \!\!\left(B\setminus\{b\}\right)$ for every $b$ in $B$.
Theorem 1.5 (Colorful Reay’s decomposition).
Fix $k\in[d].$ Let $d+k$ sets $S_{1},\dots,S_{d+k}\subset{\mathbb{R}}^{d}$ be such that for every $i\in[d+k],$ $\mathrm{pos}\ \!S_{i}$ contains a $k$-dimensional linear subspace. Then there is a rainbow sub-selection $\mathcal{R}$ from the sets $S_{1},\dots,S_{d+k}$ and its partition $\mathcal{R}=R_{1}\cup\dots\cup R_{m}$ such that
  1. (1)
    $\lvert R_{i}\rvert\geq\lvert R_{i+1}\rvert\geq 2$ for every $i\in[m-1]$.
  2. (2)
    For every $i\in[m],$ $\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{i}\right\}$ is a linear subspace of ${\mathbb{R}}^{d}$ of dimension $\sum\limits_{j\in[i]}\lvert R_{j}\rvert-i,$ and $\bigcup\limits_{j\in[i]}R_{j}$ is its positive basis.
  3. (3)
    $\mathrm{pos}\ \!\mathcal{R}=\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m}\right\}$ contains a $k$-dimensional linear subspace.
The number (of colors) $d+k$ in Theorem 1.5 is optimal as follows from the next lemma.
Lemma 1.6.
For every $d\geq k\geq 1,$ there are d+k1d+k-1 sets $S_{1},\dots,S_{d+k-1}\subset{\mathbb{R}}^{d}$ such that for every $i\in[d+k-1],$ $\mathrm{pos}\ \!S_{i}$ contains a $k$-dimensional linear subspace, but the positive cone of any rainbow sub-selection from $S_{1},\dots,S_{d+k-1}$ does not contain a $k$-dimensional subspace.
We believe that a weaker version of Theorem 1.5 with the same bound $d+k$ on the number of sets can be derived from the Colorful Carathéodory theorem (Proposition 1.4). However, we will use the fact that the size of $R_{i}$ is a monotonically decreasing function of $i$ to prove Theorem 1.3.
The following linear algebra consequence of Theorem 1.5 is simply a dual statement to Theorem 1.3, and we believe it might be of interest.
Let $A$ be a subset of ${\mathbb{R}}^{d}.$ The lineality space of $\mathrm{pos}\ \!A$ is the set $\mathrm{pos}\ \!A\cap(-\mathrm{pos}\ \!A)$, which is a linear subspace of $\mathbb{R}^{d}$. It is the unique maximal-dimensional subspace that $\mathrm{pos}\ \!A$ contains. We use $\mathrm{lpos}\ \!A$ to denote the lineality space of $\mathrm{pos}\ \!A.$
Set $h(k,d)=\max\{d+1,2(k+1)\}$ and $m(k,d)=\max\{d+1,2(d-k+1)\}.$ Note that h(k,d)=m(dk,d).h(k,d)=m(d-k,d).
Theorem 1.7.
Fix $k\in[d-1].$ Assume $A_{1},\dots,A_{d+k+1}$ are subsets of ${\mathbb{R}}^{d},$ and $\dim\mathrm{lpos}\ \!R\leq k$ for every rainbow sub-selection $R$ from these sets of size at most $h(k,d)$. Then for some index $i\in[d+k+1],$ $\dim\mathrm{lpos}\ \!A_{i}\leq k.$
The author believes that there are deep interconnections between the geometry of positive bases and Quantitative Helly-type and Carathéodory-type results, in particular with the quantitative version of Steinitz’s theorem [Ste13]. The latter asserts that if the origin in ${\mathbb{R}}^{d}$ is contained in the interior of the convex hull of a set $S$, then there exist at most $2d$ points of $S$ whose convex hull still contains the origin in its interior. Equivalently, Steinitz’s theorem implies that the cardinality of a positive basis of ${\mathbb{R}}^{d}$ is at most $2d$. The author believes that the reformulation of the Quantitative Steinitz problem in the language of positive bases is missing in the literature:
Quantitative Steinitz Problem for positive bases. Determine the largest number $r_{d}$ with the following property: whenever the Euclidean unit ball in ${\mathbb{R}}^{d}$ is contained in the convex hull of a set $S$, the set $S$ contains a subset $\{b_{1},\dots,b_{k}\}$ with $k\leq 2d$ such that
$\{b_{1}-ru,\dots,b_{k}-ru\}\quad\text{contains a positive basis of }{\mathbb{R}}^{d}$
for every unit vector $u$ and every $r\in[0,r_{d}).$
Moreover, this point of view was the main motivation for the currently best bound in the quantitative Steinitz problem [IN24, IN25, Iva25]. As suggested in [IN22] and further developed in [AHAK22], the same phenomenon also appears in quantitative Helly-type theorems via polar duality.
The rest of the paper is organized as follows: In the next section, we establish our notation and introduce several key technical lemmas that will be utilized throughout the paper. Next, we will derive Theorem 1.7 and Theorem 1.3 from Theorem 1.5 in Section 3. In Section 4, we will obtain Theorem 1.5. It is the longest and most technical proof in the paper; it does not require any other results from the paper except for the well-known results discussed in the next section. In Section 5, we will prove Lemma 1.6 and show that it implies that the numbers of colors d+(dk)+1d+(d-k)+1 and $d+k+1$ in Theorem 1.3 and Theorem 1.7 are optimal. In the final section, we will obtain a weak nonhomogeneous version of Theorem 1.3.
2. Notation, general properties of positive bases and cones
2.1. Notation
We use $[n]$ to denote the set $\{1,\dots,n\}$ for a natural $n.$ The inner product of two vectors $x$ and $y$ of ${\mathbb{R}}^{d}$ is denoted by $\left\langle x,y\right\rangle.$ The linear hull of a set $S$ is denoted by $\mathrm{Lin}\ \!{S}.$
Recall that by a rainbow sub-selection from several sets (colors), we are referring to choosing at most one element from each set (color). By a rainbow selection from several sets, we understand a choice of exactly one element from each set (color).
Recall that $h(k,d)=\max\{d+1,2(k+1)\}$ and $m(k,d)=\max\{d+1,2(d-k+1)\}.$
We recall that the polar of a set $S\subset{\mathbb{R}}^{d}$ is defined by
${S}^{\circ}=\left\{x\in{\mathbb{R}}^{d}:\;\left\langle x,s\right\rangle\leq 1\quad\text{for all}\quad s\in S\right\}.$
The set $S^{\circ\circ}$ is called the bipolar of $S.$ Recall the bipolar theorem ([PB07, Theorem 1.12.1]):
Proposition 2.1 (Bipolar theorem).
For any set $K\subset{\mathbb{R}}^{d},$ the bipolar $K^{\circ\circ}$ coincides with the closure of $\mathrm{conv}\!\left(K\cup\{0\}\right).$
2.2. Positive basis and minimal positive basis
Let $S$ be a subset of a linear space. Its positive hull, denoted $\mathrm{pos}\ \!S,$ is the set of all finite linear combinations of elements of $S$ with non-negative coefficients. We say that a set $B$ is a positive basis of a finite-dimensional linear space $L$ if $L=\mathrm{pos}\ \!B,$ but $L\neq\mathrm{pos}\ \!\!\left(B\setminus\{b\}\right)$ for any $b\in B.$
Let $B$ be a positive basis of $L.$ We say that a linear subspace $L^{\prime}$ of $L$ is a positively spanned subspace with respect to $B$ if $B\cap L^{\prime}$ is a positive basis of $L^{\prime}.$ Additionally, if $\dim L^{\prime}=n$ and $\lvert B\cap L^{\prime}\rvert=n+1,$ then we say that $L^{\prime}$ is minimal and $B\cap L^{\prime}$ is the minimal positive basis of $L^{\prime}.$ Clearly, the minimal positive basis of a non-trivial space consists of at least two vectors.
We will use next lemma, proven in [Dav54, Theorem 4.1].
Proposition 2.2.
Let $B$ be a positive basis of a finite-dimensional space $L.$ Then $B$ contains a minimal positive basis of some minimal linear subspace of $L.$
Remark 2.1.
The assertion of Theorem 1.5 implies that $R_{1}$ is a minimal positive basis of its positive hull.
2.3. Convex cones
A convex cone $C$ in ${\mathbb{R}}^{d}$ is a set of the form $a+\mathrm{pos}\ \!\!{A}$ for some $A\subset{\mathbb{R}}^{d}.$ In this case, we will call $a$ an apex of $C.$ Note that, according to our definition, an apex is not necessarily unique. Here and in the rest of the paper, we do not assume that the origin is the apex of a cone! By a non-trivial cone we understand a cone that differs from $\{0\}.$
We say that a convex cone with apex at $a$ is pointed if the cone does not contain any line. Clearly, if a convex cone is pointed, then it has a unique apex.
The following lemma is elementary and follows from the separation lemma.
Lemma 2.3.
Assume $C$ is a closed convex cone in ${\mathbb{R}}^{d}.$ If $C$ is pointed, then $C^{\circ}$ is full-dimensional in ${\mathbb{R}}^{d}.$
2.4. Lineality space
The lineality space of $\mathrm{pos}\ \!A$ is denoted by $\mathrm{lpos}\ \!A$. Surprisingly, although positive bases and lineality spaces are used extensively in several areas of mathematics, their basic properties are scattered throughout the literature (often without proofs). We will continue this tradition here, as we will need only two elementary facts.
We use $\oplus$ to denote the direct sum. Let $A$ be a set in ${\mathbb{R}}^{d}$, and denote $L=\mathrm{lpos}\ \!A$ and $L^{\perp}$ its orthogonal complement. As mentioned on page 65 of [rockafellar1970convex], one can express $\mathrm{pos}\ \!A$ as the direct sum of $L$ and $L^{\perp}\cap\mathrm{pos}\ \!A$ whenever $L$ is non-trivial. We will need the following slightly more general statement, which follows from standard linear-algebraic considerations:
Proposition 2.4.
Let $L$ be a non-trivial subspace of ${\mathbb{R}}^{d}$, and let $L^{\perp}$ be its orthogonal complement. Let $P$ denote the orthogonal projection onto $L^{\perp}$. Let $A$ be a subset of ${\mathbb{R}}^{d}$ such that $\mathrm{lpos}\ \!A\supset L$. Then
$\mathrm{pos}\ \!A=L\oplus\mathrm{pos}\ \!P(A).$
In the special case $\mathrm{lpos}\ \!A=L$, the cone $\mathrm{pos}\ \!P(A)$ is pointed.
Using the previous statement, one obtains the following:
Proposition 2.5.
Let $A$ be a subset of ${\mathbb{R}}^{d}$ with a non-trivial lineality space $\mathrm{lpos}\ \!A$. Then $A$ contains a positive basis of $\mathrm{lpos}\ \!A$.
3. Proofs of Theorem 1.7 and Theorem 1.3
We start by showing how to derive Theorem 1.7 from Theorem 1.5.
We will use the following simple arithmetic lemma. Recall that $h(k,d)=\max\{d+1,2(k+1)\}.$
Lemma 3.1.
Let $j,k,d$ be integers satisfying $1\leq j-1\leq k\leq d-1.$ Then $\frac{jk}{j-1}+j\leq h(k,d).$
Proof.
The lemma was proved in [Bár24, Claim 3.1]. For the sake of completeness, we provide an alternative short computation:
$\frac{jk}{j-1}+j\leq 2(k+1)\quad\Longleftrightarrow\quad j-2\leq\frac{j-2}{j-1}k.$
The latter inequality holds whenever $1\leq j-1\leq k$, which is exactly our assumption. The lemma follows since $2(k+1)\leq h(k,d).$
Let us continue with the proof of Theorem 1.7. We essentially copy the proof given in [Bár24] for the monochromatic version of the result.
Proof of Theorem 1.7.
Define $L_{i}=\mathrm{lpos}\ \!A_{i},$ $i\in[d+k+1],$ and $t=\min\limits_{i\in[d+k+1]}\dim L_{i}.$ Assume that $t\geq k+1,$ otherwise, there is nothing to prove. By Theorem 1.5, there is a rainbow sub-selection $\mathcal{R}$ from the sets $A_{1},\dots,A_{d+k+1}$, together with its partition $\mathcal{R}=R_{1}\cup\cdots\cup R_{m}$ such that
  1. (1)
    $\lvert R_{i}\rvert\geq\lvert R_{i+1}\rvert\geq 2$ for every $i\in[m-1]$.
  2. (2)
    For every $i\in[m],$ $\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{i}\right\}$ is a linear subspace of ${\mathbb{R}}^{d}$ of dimension $\sum\limits_{j\in[i]}\lvert R_{j}\rvert-i,$ and $\bigcup\limits_{j\in[i]}R_{j}$ is its positive basis.
  3. (3)
    $\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m}\right\}$ contains a $(k+1)$-dimensional linear subspace.
The last property means that $R_{1}\cup\dots\cup R_{m}$ is a positive basis of at least $(k+1)$-dimensional linear space $\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m}\right\}.$
We now prove by induction on $j$ that $\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\ldots\cup R_{j}\right\}\leq k.$
The base case $j=1.$ Since $R_{1}$ is a subset of ${\mathbb{R}}^{d},$ $\dim\mathrm{pos}\ \!R_{1}\leq d.$ As $R_{1}$ is a minimal positive basis of $\mathrm{pos}\ \!R_{1},$ we have $\lvert R_{1}\rvert\leq d+1\leq h(k,d).$ By the assumption of the theorem, $\dim{\mathrm{lpos}\ \!R_{1}}\leq k.$
For the inductive step $j-1\to j$, assume that $j\geq 2.$ Since $R_{j}$ has the smallest size among the sets $R_{1},\ldots,R_{j},$ we have
(1) $\lvert R_{j}\rvert\leq\frac{1}{j-1}\sum\limits_{i\in[j-1]}\lvert R_{i}\rvert$.
By the choice of the sets $R_{1},\dots,R_{j-1}$ and by the induction hypothesis,
(2) $\sum\limits_{i\in[j-1]}\lvert R_{i}\rvert=\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{j-1}\right\}+(j-1)\leq k+(j-1).$
Since each $\lvert R_{i}\rvert\geq 2$, we have $\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\dots\cup R_{j}\right\}\geq j$ for every $j\in[m]$. Therefore,
$j\leq\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{j}\right\}=\sum\limits_{i\in[j-1]}\lvert R_{i}\rvert+\lvert R_{j}\rvert-j$
[…]
Consequently, $j\leq\frac{jk}{j-1},$ which implies that $j-1\leq k.$
We can now apply Lemma 3.1. It implies that
$\sum\limits_{i\in[j]}\lvert R_{j}\rvert=\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{j}\right\}+j\leq h(k,d).$
Thus, by the assumption of the theorem, $\dim\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{j}\right\}\leq k.$
Finally, for $j=m,$ we get $k+1\leq\dim\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m}\right\}\leq k.$ This contradiction shows that our assumption was false. We conclude $t\leq k,$ completing the proof of Theorem 1.7. ∎
To prove Theorem 1.3, we need the following simple technical lemma.
Lemma 3.2.
Assume $A$ is a finite set of nonzero vectors in ${\mathbb{R}}^{d}.$ We use $L^{\perp}$ to denote the orthogonal complement of the lineality space $\mathrm{lpos}\ \!A,$ and $P$ to denote the orthogonal projection onto $L^{\perp}.$ Denote the solution set of the homogeneous system $\left\langle a,x\right\rangle\leq 0,a\in A$ by $S.$ Then $S$ is a full-dimensional cone within $L^{\perp},$ and the identities
$S={\left(\mathrm{pos}\ \!A\right)}^{\circ}=L^{\perp}\cap{\left(\mathrm{pos}\ \!P(A)\right)}^{\circ}$
hold.
Proof.
The leftmost identity $S={\left(\mathrm{pos}\ \!A\right)}^{\circ}$ immediately follows from the definition of the polar set.
By Proposition 2.4, $\mathrm{pos}\ \!A=\mathrm{lpos}\ \!A\oplus\mathrm{pos}\ \!P(A),$ and $\mathrm{pos}\ \!P(A)$ is pointed within ${\mathbb{R}}^{d},$ and consequently, within $L^{\perp}$. Thus, by the standard properties of the polar transform ( see, for example, [Sch14, Theorem 1.6.9]),
${\left(\mathrm{pos}\ \!A\right)}^{\circ}={\left(\mathrm{lpos}\ \!A\oplus\mathrm{pos}\ \!P(A)\right)}^{\circ}=L^{\perp}\cap{\left(\mathrm{pos}\ \!P(A)\right)}^{\circ}.$
Since the cone $\mathrm{pos}\ \!P(A)$ is a pointed cone within $L^{\perp},$ Lemma 2.3 implies that the cone $L^{\perp}\cap{\left(\mathrm{pos}\ \!P(A)\right)}^{\circ}$ is full-dimensional within $L^{\perp}.$
Proof of Theorem 1.3.
By Lemma 3.2, the system $\left\langle a,x\right\rangle\leq 0,a\in\mathcal{A}_{i}$ has at least $k$ linearly independent solutions if and only if the dimension of $\left(\mathrm{lpos}\ \!\mathcal{A}_{i}\right)^{\perp}$ is at least $k.$ Or equivalently, the dimension of $\mathrm{lpos}\ \!\mathcal{A}_{i}$ is at most dk.d-k. Similarly, for a rainbow sub-selection $R$ from the sets $\mathcal{A}_{1},\dots,\mathcal{A}_{d+(d-k)+1},$ the system $\left\langle a,x\right\rangle\leq 0,\;a\in R$ has at least $k$ linearly independent solutions if and only if the dimension of $\mathrm{lpos}\ \!R$ is at most dk.d-k. The desired implication follows from Theorem 1.7. ∎
4. Proof of Theorem 1.5
We start by reducing the problem to an easier one.
Theorem 4.1.
Fix $k\in[d].$ Let $d+k$ sets $S_{1},\dots,S_{d+k}\subset{\mathbb{R}}^{d}$ be such that for every $i\in[d+k],$ $\mathrm{pos}\ \!S_{i}$ contains a $k$-dimensional linear subspace. Then there is a rainbow sub-selection $R_{1}\cup\dots\cup R_{m}$ from the sets such that
  1. (1)
    $\lvert R_{i}\rvert\geq\lvert R_{i+1}\rvert\geq 2$ for all $i\in[m-1]$.
  2. (2)
    For every $i\in[m],$ $\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{i}\right\}$ is a linear subspace of ${\mathbb{R}}^{d}$ of dimension $\sum\limits_{j\in[i]}\lvert R_{j}\rvert-i.$
  3. (3)
    $\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m}\right\}$ contains a $k$-dimensional linear subspace.
The only difference between Theorem 1.5 and Theorem 4.1 is that in the latter we do not state that $\bigcup\limits_{j\in[i]}R_{j}$ is a positive basis of $\mathrm{pos}\ \!\!\left\{R_{1}\cup\cdots\cup R_{i}\right\},$ $i\in[m].$ However, it is easy to see that Theorem 4.1 implies Theorem 1.5.
Proof of Theorem 1.5.
By Theorem 4.1, there is a rainbow sub-selection $R^{\prime}$ from $S_{1},\dots,S_{d+k}$ such that $\mathrm{pos}\ \!R^{\prime}$ is a linear subspace and $\dim\mathrm{pos}\ \!R^{\prime}\geq k.$ By Proposition 2.5, $R^{\prime}$ contains a positive basis, say $R,$ of $\mathrm{pos}\ \!R^{\prime}.$ Applying Theorem 4.1 to $d+\dim\mathrm{pos}\ \!R^{\prime}$ copies of $R,$ we get the desired set system. ∎
From now on, we proceed with the proof of Theorem 4.1.
Our proof is based on an inductive construction. First, we show how to choose the first rainbow set $R_{1},$ then we will show how to find a next rainbow set $R_{m}$ provided that the dimension of $\mathrm{lpos}\ \!\!\left\{R_{1}\cup\dots\cup R_{m-1}\right\}$ is strictly less than $k.$ Finally, we will show that the process terminates after a finite number of steps and the constructed sets satisfy the desired properties.
4.1. Choice of the first rainbow set
We start with the simplest case of $k=1.$
Lemma 4.2.
Assume $L$ is an $n$-dimensional vector space. Assume that for every $i\in[n+1],$ a finite set $B_{i}$ is a positive basis of a non-trivial subspace $H_{i}$ of $L.$ Then there is a rainbow selection $R$ from the sets $B_{1},\dots,B_{n+1}$ such that $\mathrm{lpos}\ \!R$ is a non-trivial subspace of $L.$
Proof.
Fix $i\in[n+1].$ Since $B_{i}$ is a positive basis of $H_{i},$ $0\in\mathrm{conv}{B_{i}}.$ Thus, by the colorful Carathéodory theorem – Proposition 1.4, there is a rainbow selection $R$ such that $0\in\mathrm{conv}{R}.$ Since the original sets are positive bases in the corresponding subspaces, they do not contain the origin. Thus, the origin belongs to the relative interior of a face $F$ of $\mathrm{conv}{R}$ satisfying $\dim F\geq 1.$ It means that $\mathrm{lpos}\ \!R\supset\mathrm{Lin}\ \!{F}.$ The lemma follows. ∎
Using Proposition 2.5, Proposition 2.2, and Lemma 4.2 together, we get
Corollary 4.3.
Assume $L$ is an $n$-dimensional vector space. Assume that for every $i\in[n+1],$ a set $S_{i}$ satisfies $\dim\mathrm{lpos}\ \!S_{i}\geq 1.$ Then there is a non-empty rainbow sub-selection $R$ from the sets $S_{1},\dots,S_{n+1}$ such that $R$ is a minimal basis of the minimal subspace $\mathrm{pos}\ \!R.$
By Corollary 4.3, there is a non-empty subset $I$ of $[d+k]$ and a rainbow sub-selection $R$ from $\bigcup\limits_{i\in I}S_{i}$ such that $R$ is a minimal positive basis of the minimal subspace $\mathrm{pos}\ \!R.$ Let $I_{1}\subset[d+k]$ be an index set of maximal cardinality for which there is a rainbow sub-selection $R_{1}$ from the union $\bigcup\limits_{i\in I_{1}}S_{i}$ such that $R_{1}$ is a minimal positive basis of the minimal subspace $\mathrm{pos}\ \!R_{1}.$
Note that automatically:
(3) $\lvert R_{1}\rvert\geq 2,\ \mathrm{pos}\ \!R_{1}=\mathrm{lpos}\ \!R_{1},\quad\text{and}\quad\dim\mathrm{pos}\ \!R_{1}=\lvert R_{1}\rvert-1.$
4.2. Inductive construction
Assume that for some $m^{\prime}>1,$ the sets $R_{1},\dots,R_{m^{\prime}-1}$ have already been chosen.
If the dimension of $\mathrm{lpos}\ \!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}$ is at least $k,$ we terminate the construction process.
If the dimension of $\mathrm{lpos}\ \!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}$ is strictly less than $k,$ we proceed as follows.
Denote $\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}$ by $L_{m^{\prime}-1}.$ We use $C_{m^{\prime}-1}$ to denote the orthogonal complement of $L_{m^{\prime}-1},$ and $P_{m^{\prime}-1}$ to denote the orthogonal projection onto $C_{m^{\prime}-1}.$ We set $I_{m^{\prime}}\subset[k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ to be a non-empty index set of maximal cardinality such that there is a rainbow sub-selection $R_{m^{\prime}}$ from the sets $S_{i},i\in I_{m^{\prime}},$ such that $P_{m^{\prime}-1}\ \!\!\!\left(R_{m^{\prime}}\right)$ is a minimal positive basis of the minimal linear subspace $\mathrm{pos}\ \!P_{m^{\prime}-1}\ \!\!\!\left(R_{m^{\prime}}\right).$
4.3. Correctness of the inductive construction
We divide the proof into several simple steps.
Now we show that it is possible to add a new rainbow set if the dimension of the lineality space of the union of already chosen ones is strictly less than $k.$
Lemma 4.4.
Assume that for some $m^{\prime}>1,$ the sets $R_{1},\dots,R_{m^{\prime}-1}$ have already been chosen in such a way that
  1. (1)
    $\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}<k.$
  2. (2)
    $\lvert R_{i}\rvert\geq 2$ for each $i\in[m^{\prime}-1]$.
  3. (3)
    For every $i\in[m^{\prime}-1],$ $\dim\mathrm{lpos}\ \!\!\!\left(R_{1}\cup\cdots\cup R_{i}\right)=\sum\limits_{j\in[i]}\lvert R_{j}\rvert-i.$
Then there is an index set $I\subset[k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ and a rainbow sub-selection $R$ from the sets $S_{i},i\in I,$ such that $P_{m^{\prime}-1}\ \!\!\!\left(R\right)$ is a minimal positive basis of the minimal linear subspace $\mathrm{pos}\ \!P_{m^{\prime}-1}\ \!\!\!\left(R\right).$
Proof.
For any $i,$ the dimension of the space $\mathrm{lpos}\ \!P_{m^{\prime}-1}(\mathrm{lpos}\ \!S_{i})$ is at least $\dim\mathrm{lpos}\ \!S_{i}-\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}\geq 1.$ That is, $\mathrm{lpos}\ \!P_{m^{\prime}-1}(\mathrm{lpos}\ \!S_{i})$ is a non-trivial subspace of $C_{m^{\prime}-1}.$ On the other hand, the cardinality of the set $k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ equals
$d+k-\sum\limits_{j\in[m^{\prime}-1]}\lvert R_{j}\rvert=d-\left(\sum\limits_{j\in[m^{\prime}-1]}\lvert R_{j}\rvert-(m^{\prime}-1)\right)+(k-(m^{\prime}-1)).$
By our assumption, the cardinality of the set $k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ equals
(4) $d-\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}+(k-(m^{\prime}-1)).$
Since $\lvert R_{i}\rvert\geq 2$ for all $i\in[m^{\prime}-1],$ we get
$k>\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}=\sum\limits_{j\in[m^{\prime}-1]}\lvert R_{j}\rvert-(m^{\prime}-1)\geq m^{\prime}-1.$
Hence, $k-(m^{\prime}-1)\geq 1.$ However, $\dim C_{m^{\prime}-1}=d-\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}.$ Thus, using (4) in Corollary 4.3, we conclude that there is a non-empty subset $I$ of $k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ and a rainbow sub-selection $R^{\prime}$ from the sets $P_{m^{\prime}-1}(S_{i}),i\in I,$ such that $R^{\prime}$ is a minimal positive basis of the minimal linear subspace $\mathrm{lpos}\ \!R^{\prime}.$ There are $s_{i}\in S_{i},$ $i\in I$ such that $R^{\prime}=\bigcup\limits_{i\in I}P_{m^{\prime}-1}(s_{i}).$ We complete the proof by taking $R=\bigcup\limits_{i\in I}\left\{s_{i}\right\}.$
Corollary 4.5.
Assume that for some $m^{\prime}>1,$ the sets $R_{1},\dots,R_{m^{\prime}-1}$ have already been chosen by our inductive construction in such a way that $\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}<k.$ Then there are $I_{m^{\prime}}\subset[k+d]\setminus\left(I_{1}\cup\dots\cup I_{m^{\prime}-1}\right)$ of maximal cardinality and a rainbow sub-selection $R_{m^{\prime}}$ from $S_{i},i\in I_{m^{\prime}},$ such that $P_{m^{\prime}-1}\ \!\!\!\left(R_{m^{\prime}}\right)$ is a minimal positive basis of the minimal linear subspace $\mathrm{pos}\ \!P_{m^{\prime}-1}\ \!\!\!\left(R_{m^{\prime}}\right).$
Now let us show that the other desired properties of $R_{m^{\prime}}$ hold.
We will need the following lemma, which is essentially Lemma 2.5 of [Rea65].
Lemma 4.6.
Let $S$ be a subset of a linear space. Assume
  1. (1)
    There is a subset $R$ of $S$ of maximal cardinality such that $R$ is a minimal positive basis of the minimal subspace $\mathrm{pos}\ \!R.$
  2. (2)
[…]
Then $\lvert Q\rvert\leq\lvert R\rvert$, $\mathrm{pos}\ \!\!\!\left\{R\cup Q\right\}=\mathrm{lpos}\ \!\!\!\left\{R\cup Q\right\}=\mathrm{pos}\ \!\!\!\left\{R\cup P(Q)\right\},$ and $\dim\mathrm{pos}\ \!\!\!\left\{R\cup Q\right\}=\lvert Q\rvert+\lvert R\rvert-2.$
Proof.
[…]
Assume $\sum_{j\in J}\beta_{j}q_{j}=0.$ Then the origin is in the relative interior of $\mathrm{conv}\ \!Q,$ and therefore $\mathrm{pos}\ \!\ \!Q$ is a linear subspace of dimension $\lvert Q\rvert-1.$ In this case the choice of $R$ implies the inequality $\lvert R\rvert\geq\lvert Q\rvert$. Moreover, in this case, the subspace $\mathrm{pos}\ \!\left\{R\cup Q\right\}$ is equal to the direct sum of the linear subspaces $\mathrm{pos}\ \!R$ and $\mathrm{pos}\ \!Q.$ Hence, $\dim\mathrm{pos}\ \!\!\left\{R\cup Q\right\}=\lvert R\rvert+\lvert Q\rvert-2$ and $R\cup Q$ is a positive basis of $\mathrm{pos}\ \!\!\left\{R\cup Q\right\}.$
Assume now that $\sum_{j\in J}\beta_{j}q_{j}\neq 0.$ Since $R$ is a minimal basis of $\mathrm{pos}\ \!R,$ there is a unique proper subset $I^{\prime}$ of $I$ of minimal cardinality and positive coefficients $\alpha_{i},i\in I^{\prime},$ such that
$\sum_{j\in J}\beta_{j}q_{j}+\sum_{i\in I^{\prime}}\alpha_{i}r_{i}=0.$
This equation and the minimality of $\lvert I^{\prime}\rvert$ imply that the origin is in the relative interior of $\mathrm{conv}\left(R^{\prime}\cup Q\right),$ where $R^{\prime}=\{r_{i}:\;i\in I^{\prime}\}.$ Hence, $\mathrm{pos}\ \!\!\!\left\{R^{\prime}\cup Q\right\}=\mathrm{Lin}\ \!{\ \!\!\!\left\{R^{\prime}\cup Q\right\}}.$ However, the construction shows that $\dim\left(R^{\prime}\cup Q\right)=\lvert I^{\prime}\rvert+\left(\lvert Q\rvert-1\right).$ Hence $R^{\prime}\cup Q$ is the minimal positive basis for $\mathrm{pos}\ \!\!\left\{R^{\prime}\cup Q\right\}.$ By the choice of $R,$ $\lvert R\rvert\geq\lvert Q\cup R^{\prime}\rvert>\lvert Q\rvert$.
Since $\mathrm{pos}\ \!R=\mathrm{lpos}\ \!R$ by the choice of $R,$ we have
$\mathrm{pos}\ \!\!\left\{R\cup Q\right\}=\mathrm{pos}\ \!\!\left\{R\cup P(Q)\right\}.$
Clearly, $\lvert P(Q)\rvert=\lvert Q\rvert$. However, $\mathrm{pos}\ \!\!\!\left\{R\cup P(Q)\right\}$ is the direct sum of $\mathrm{pos}\ \!R$ and $\mathrm{pos}\ \!P(Q),$ which is a linear subspace of dimension $\lvert R\rvert+\lvert Q\rvert-2$ by the choice of $R$ and $Q.$ The proof of Lemma 4.6 is complete. ∎
Lemma 4.7.
Assume that for some $m^{\prime}>1,$ the sets $R_{1},\dots,R_{m^{\prime}-1}$ have already been chosen by our inductive construction in such a way that
  1. (1)
    $\dim\mathrm{lpos}\ \!\!\!\left\{R_{1}\cup\dots\cup R_{m^{\prime}-1}\right\}<k.$
  2. (2)
    $\lvert R_{i}\rvert\geq 2$ for each $i\in[m^{\prime}-1]$.
  3. (3)
    For every $i\in[m^{\prime}-1],$ $\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{i}\right\}$ is a linear subspace of ${\mathbb{R}}^{d}$ of dimension $\sum\limits_{j\in[i]}\lvert R_{j}\rvert-i.$
Then the inductive construction described in Subsection 4.2 returns $R_{m^{\prime}}$ satisfying the following properties:
  1. (1)
    $\lvert R_{m^{\prime}}\rvert\geq 2.$
  2. (2)
    $\lvert R_{m^{\prime}-1}\rvert\geq\lvert R_{m^{\prime}}\rvert$.
  3. (3)
    $\mathrm{pos}\ \!\!\!\left\{R_{1}\cup\cdots\cup R_{m^{\prime}}\right\}$ is a linear subspace of ${\mathbb{R}}^{d}$ of dimension $\sum\limits_{j\in[m^{\prime}]}\lvert R_{j}\rvert-m^{\prime}.$
Proof.
By Corollary 4.5, the inductive construction described in Subsection 4.2 returns $R_{m^{\prime}}.$ By construction, $\lvert R_{m^{\prime}}\rvert\geq 2.$
We use $P_{0}$ to denote the identity operator on ${\mathbb{R}}^{d},$ $L_{0}$ and $C_{0}$ to denote 0 and ${\mathbb{R}}^{d},$ respectively. Set $R=P_{m^{\prime}-2}(R_{m^{\prime}-1})$ and $Q=P_{m^{\prime}-2}(R_{m^{\prime}}).$ Since $C_{m^{\prime}-2}\supset C_{m^{\prime}-1}\supset C_{m^{\prime}},$
$P_{m^{\prime}-1}(Q)=P_{m^{\prime}-1}P_{m^{\prime}-2}(R_{m^{\prime}})=P_{m^{\prime}-1}(R_{m^{\prime}}).$
Hence, $\lvert R\rvert=\lvert R_{m^{\prime}-1}\rvert$ and $\lvert Q\rvert=\lvert R_{m^{\prime}}\rvert$. Moreover, $R$ is a minimal basis of $\mathrm{pos}\ \!R$ by the choice of $R_{m^{\prime}-1}.$ Thus, the sets $R$ and $Q$ satisfy the assumptions of Lemma 4.6 with $S=R\cup Q$. Consequently, $\lvert R_{m^{\prime}-1}\rvert\geq\lvert R_{m^{\prime}}\rvert$ and $\mathrm{pos}\ \!\!\!\left\{R\cup Q\right\}$ is a linear subspace of dimension $\lvert R_{m^{\prime}-1}\rvert+\lvert R_{m^{\prime}}\rvert-2.$
The lemma follows in the case $m^{\prime}=2.$ If $m^{\prime}>2,$ $L_{m^{\prime}-2}$ is non-trivial. By Proposition 2.4, applied with $L=L_{m^{\prime}-2}$ and $P=P_{m^{\prime}-2}$, we obtain
$\mathrm{pos}\ \!L_{m^{\prime}}=\mathrm{pos}\ \!\!\bigl(L_{m^{\prime}-2}\cup R_{m^{\prime}-1}\cup R_{m^{\prime}}\bigr)=\mathrm{pos}\ \!L_{m^{\prime}-2}\;\oplus\;\mathrm{pos}\ \!\!\bigl(P_{m^{\prime}-2}(R_{m^{\prime}-1})\cup P_{m^{\prime}-2}(R_{m^{\prime}})\bigr).$
Since $R=P_{m^{\prime}-2}(R_{m^{\prime}-1})$ and $Q=P_{m^{\prime}-2}(R_{m^{\prime}})$, the above becomes
$\mathrm{pos}\ \!L_{m^{\prime}}=\mathrm{pos}\ \!L_{m^{\prime}-2}\;\oplus\;\mathrm{pos}\ \!\!\left\{R\cup Q\right\}.$
Thus,
$\dim\mathrm{pos}\ \!L_{m^{\prime}}=\dim\mathrm{pos}\ \!L_{m^{\prime}-2}+\bigl(\lvert R_{m^{\prime}-1}\rvert+\lvert R_{m^{\prime}}\rvert-2\bigr).$
Using the induction hypothesis,
$\dim\mathrm{pos}\ \!L_{m^{\prime}-2}=\sum_{j\in[m^{\prime}-2]}\lvert R_{j}\rvert-(m^{\prime}-2),$
we obtain
$\dim\mathrm{pos}\ \!L_{m^{\prime}}=\sum_{j\in[m^{\prime}-2]}\lvert R_{j}\rvert-(m^{\prime}-2)+\lvert R_{m^{\prime}-1}\rvert+\lvert R_{m^{\prime}}\rvert-2=\sum_{j\in[m^{\prime}]}\lvert R_{j}\rvert-m^{\prime}.$
Let us combine together all the small lemmas to prove Theorem 4.1.
Proof of Theorem 4.1.
Choose $R_{1}$ as described in Subsection 4.1. The existence of such a choice was justified in Subsection 4.1.
If $\lvert R_{1}\rvert>k,$ then the inductive construction terminates immediately at $m=1$ and by (3) $R_{1}$ is the desired rainbow sub-selection.
If $\lvert R_{1}\rvert\leq k,$ we use the inductive construction from Subsection 4.2 to construct sets $R_{2},\dots,R_{m}$ while $\dim\mathrm{lpos}\ \!\!\left\{R_{1}\cup\dots\cup R_{m-1}\right\}<k.$ By Corollary 4.5 and (3), we can choose the next rainbow set as described in the inductive construction. By Lemma 4.7, the constructed set system satisfies the desired properties. Since we add at least one dimension at each step, the process terminates after at most $k$ steps. ∎
5. Optimality of Helly numbers
We start by proving Lemma 1.6, which shows the optimality of $d+k$ in Theorem 1.5.
Proof of Lemma 1.6.
We proceed by induction on $d$.
For $d=k$, set
$S_{1}=\dots=S_{2k-1}=\left\{\pm e_{1},\dots,\pm e_{k}\right\}.$
Any rainbow sub-selection $R$ from these sets omits at least one of the $2k$ points $\left\{\pm e_{1},\dots,\pm e_{k}\right\}$. Hence, $\mathrm{pos}\ \!R$ cannot contain the whole space ${\mathbb{R}}^{k}$.
Now we explain how to add one dimension and one new set to our set system. We assume that ${\mathbb{R}}^{k}$ is the subspace of ${\mathbb{R}}^{d}$ spanned by its first $k$ standard basis vectors $e_{1},\dots,e_{k}$ for any $k\in[d]$.
The inductive construction is as follows: assume that for some $d>k$ we have already constructed sets
$S_{1},\dots,S_{d+k-2}\subset{\mathbb{R}}^{d-1}.$
We then define $S_{d+k-1}$ to be any positive basis of ${\mathbb{R}}^{d}$ whose vectors do not lie in ${\mathbb{R}}^{d-1}$.
We now show that this construction indeed produces the desired sets. More precisely, we will show that the lineality space of any rainbow sub-selection $R$ coincides with the lineality space of the elements of $R$ taken from the first 2k12k-1 sets, that is,
$\mathrm{lpos}\ \!R=\mathrm{lpos}\ \!\!\left(R\cap\left(S_{1}\cup\dots\cup S_{2k-1}\right)\right).$
Since $\mathrm{pos}\ \!S_{1}=\dots=\mathrm{pos}\ \!S_{2k-1}={\mathbb{R}}^{k}$ and $\mathrm{pos}\ \!S_{2k-1+i}={\mathbb{R}}^{k+i}$, and since we have already settled the case $d=k$, this identity of lineality spaces implies the desired bound
$\dim\mathrm{lpos}\ \!R<k.$
Assume now that $\mathrm{lpos}\ \!R$ is non-trivial; otherwise, there is nothing to prove. By Proposition 2.5, the set $R$ contains a positive basis $F$ of its lineality space. Let $j$ be the largest index such that $S_{j}\cap F$ is non-empty. If $j\leq 2k-1$, there is again nothing to prove. If j>2k1j>2k-1, then the single element of $S_{j}\cap F$ has a non-zero (j(k1))(j-(k-1))-th coordinate, and it is the only element of $F$ whose (j(k1))(j-(k-1))-th coordinate is non-zero. This contradicts the fact that $F$ is a positive basis of a non-trivial subspace.
The proof of Lemma 1.6 is complete. ∎
Remark 5.1.
The sets $S_{i},i>2k-1,$ can be chosen in such a way that $\dim\mathrm{lpos}\ \!S_{i}=k$.
Corollary 5.1.
The number of colors $d+k+1$ in Theorem 1.7 is optimal. For any $k\in[d-1],$ there are subsets $A_{1},\dots,A_{d+k}$ of ${\mathbb{R}}^{d}$ with $\dim\mathrm{lpos}\ \!A_{i}>k,$ $i\in[d+k],$ and such that $\dim\mathrm{lpos}\ \!R\leq k$ for every rainbow sub-selection $R$ from these sets.
Remark 5.2.
We do not assume that the size of $R$ is at most $h(k,d)$.
Proof.
Since $k\in[d-1],$ $1<k+1\leq d$. Thus, any $d+k$ sets returned by Lemma 1.6 for $k\to k+1$ provide the desired example. ∎
Similarly, one gets
Corollary 5.2.
The number of colors d+(dk)+1d+(d-k)+1 in Theorem 1.3 is optimal. For any $k\in[d-1],$ there are $\mathcal{A}_{1},\dots,\mathcal{A}_{d+(d-k)}$ finite subsets of nonzero vectors in ${\mathbb{R}}^{d}$ such that the dimension of the solution set of the systems $\left\langle a,x\right\rangle\leq 0,\;a\in\mathcal{A}_{i},$ $i\in[d+(d-k)]$ is strictly less than $k$, but for every rainbow sub-selection $R$ from these sets the system $\left\langle a,x\right\rangle\leq 0,\;a\in R$ has at least $k$ linearly independent solutions.
Proof.
Since $k\in[d-1],$ $1<d-k+1\leq d$. By Lemma 3.2, any d+(dk)d+(d-k) sets returned by Corollary 5.1 for $k\to d-k$ provide the desired example. ∎
As mentioned in the Introduction, the optimality of $h(k,d)$ and $m(k,d)$ was shown in [Bár24]. For completeness, we provide the example that shows the optimality of $h(k,d)$ (the optimality of $m(k,d)$ follows immediately from it and Lemma 3.2).
Consider the case $d+1\geq 2(k+1)$ in which $h(k,d)=d+1$. Set $A$ to be the vertices $v_{1},\dots,v_{d+1}$ of a regular simplex with $v_{1}+\dots+v_{d+1}=0$. Then, $\mathrm{lpos}\ \!A={\mathbb{R}}^{d}$ and for any $B\subset A$ of size at most $d,$ $\mathrm{lpos}\ \!B=\{0\}$.
Consider the case $d+1<2(k+1)$ in which $h(k,d)=2(k+1)$. Set $A$ to be the set of vertices of the $(k+1)$-dimensional cross-polytope $\left\{\pm e_{1},\dots,\pm e_{k+1}\right\}$. Clearly, one has $\dim\mathrm{lpos}\ \!A=k+1$ but $\dim\mathrm{lpos}\ \!\left(A\setminus\left\{a\right\}\right)\leq k$ for every $a\in A$.
6. Nonhomogeneous Linear Systems
Proposition 1.2 was a byproduct of the proof method of the following result [Bár24, Theorem 1.2]:
Proposition 6.1.
Fix $k\in[d-1]$. Assume $\mathcal{F}$ is a finite family of convex sets in ${\mathbb{R}}^{d}$. If the intersection of any subfamily of at most $m(k,d)$ sets contains a $k$-dimensional cone, then the intersection $\bigcap\limits_{F\in\mathcal{F}}F$ of all sets in the family contains a $k$-dimensional cone.
Let us show that Proposition 6.1 and Proposition 1.2 are equivalent. Using a somewhat standard reduction procedure, one can equivalently assume that $\mathcal{F}$ is a finite family of the solution sets of finite nonhomogeneous systems of linear inequalities. The following trivial statement allows us to move a convex cone inside a convex set:
Lemma 6.2.
Let $K\subset{\mathbb{R}}^{d}$ be a closed convex set containing a cone $C$ with apex at $u$. Then $K$ contains the cone vu+Cv-u+C for every $\in K; vv$ is an apex of this cone.
Using the fact that $m(k,d)\geq d+1$ in the classical Helly’s theorem and Lemma 6.2, we see that one can equivalently assume that $\mathcal{F}$ is a finite family of the solution sets of finite homogeneous systems of linear inequalities. Thus, Proposition 6.1 and Proposition 1.2 are indeed equivalent.
Unfortunately, the previous argument does not apply directly in the colorful case. We can show the following weaker colorful version of Proposition 6.1, which is essentially a nonhomogeneous version of Theorem 1.3.
Theorem 6.3.
Fix $k\in[d-1]$. Assume $\mathcal{F}_{1},\dots,\mathcal{F}_{d+(d-k)+1}$ are finite families of convex sets in ${\mathbb{R}}^{d}$ such that the intersection of sets in any rainbow sub-selection $R$ of size at most $m(k,d)$ from these sets contains a $k$-dimensional cone. Then for some $i\in[d+(d-k)+1]$, all sets in $\mathcal{F}_{i}$ contain a translate of some $k$-dimensional cone $C$.
Our proof strategy is to reduce the problem to the homogeneous case obtained in Theorem 1.3.
We need the following two properties of convex cones.
Lemma 6.4.
Assume $C$ is a closed convex cone in ${\mathbb{R}}^{d}$ with apex at the origin, and $c_{1},\dots,c_{n}$ are points of $C$. Then
$c_{1}+\dots+c_{n}+C\subset\bigcap\limits_{i\in[n]}\left(c_{i}+C\right).$
Proof.
The lemma follows from Lemma 6.2 and the observation that $c_{1}+\dots+c_{n}-c_{i}\in C$ for every $i\in[n]$. ∎
Lemma 6.5.
Assume $K$ is a closed convex set in ${\mathbb{R}}^{d}$ and a non-trivial closed convex cone $C$ with apex at the origin is contained within $K^{\circ\circ}$. Then $K$ contains a translate of $C$.
Proof.
Since $K^{\circ\circ}$ contains a point different from the origin, $K$ is non-empty; say $p\in K$. Take any vector $v$ within $C$ and $t>2$. Since $tv$ belongs to the closure of $\mathrm{conv}\left(K\cup\{0\}\right)$, there is a vector $v_{t}$ in $K$ such that the segment with endpoints 0 and $v_{t}$ intersects the unit ball centered at $tv$, that is, $\lambda_{t}v_{t}=tv+\delta_{t}$ with $\left|\delta_{t}\right|\leq 1$ and $\lambda_{t}\in[0,1]$. By convexity, the point $p+\lambda(v_{t}-p)$ is in $K$ for all $\lambda\in[0,1]$. Thus,
$K\supset p+\frac{\lambda_{t}}{t}(v_{t}-p)=p\left(1-\frac{\lambda_{t}}{t}\right)+\frac{tv+\delta_{t}}{t}\to p+v\quad\text{as}\quad t\to\infty.$
Hence, $p+v\in K$, and consequently, $p+C\subset K.$
Now, we are ready to prove Theorem 6.3, a non-homogeneous version of Theorem 1.3.
Proof of Theorem 6.3.
The idea is to reduce the original families to families of closed polyhedral convex cones with apexes at the origin, that is, to the situation resolved in Theorem 1.3.
Without loss of generality, we can assume that all the sets are closed. Indeed, a convex set contains a non-trivial cone if and only if the closure of the set contains the closure of the cone. Also, we disregard the sets that coincide with the whole space ${\mathbb{R}}^{d}$.
For every rainbow sub-selection $R$ from the families $\mathcal{F}_{1},\dots,\mathcal{F}_{d+(d-k)+1}$, we use $C(R)$ to denote a $k$-dimensional cone of the form $\mathrm{pos}\ \!\!\!\left\{v_{1}(R),\dots,v_{k}(R)\right\}$ with linearly independent vectors $v_{1}(R),\dots,v_{k}(R)$ such that its translate is contained in the intersection set of $R$.
For every $i\in[d+(d-k)+1]$ and every $F_{i}\in\mathcal{F}_{i}$, define $F_{i}^{\prime}$ as the convex hull of the cones $C(R)$ for all the rainbow sub-selections $R$ of size at most $m(k,d)$ from $\mathcal{F}_{1},\dots,\mathcal{F}_{d+(d-k)+1}$ containing $F_{i}$. Since $K$ and the origin are contained within $K^{\circ\circ}$, Lemma 6.2 implies that every such $C(R)$ is contained in $F_{i}^{\circ\circ}$. Hence, $F_{i}^{\prime}\subset F_{i}^{\circ\circ}$. Also, $F_{i}^{\prime}$ is a polyhedral cone with an apex at the origin, that is, $F_{i}^{\prime}$ is the solution set of a finite homogeneous system of linear inequalities $\left\langle a_{i},x\right\rangle\leq 0$, $a_{i}\in A_{i}\subset{\mathbb{R}}^{d}\setminus\{0\}$.
Define $\mathcal{F}_{1}^{\prime},\dots,\mathcal{F}_{d+(d-k)+1}^{\prime}$ by
$\mathcal{F}_{i}^{\prime}=\left\{F_{i}^{\prime}:\;F_{i}\in\mathcal{F}_{i}\right\},\qquad i\in[d+(d-k)+1].$
On the one hand, we have that the intersection set of any rainbow sub-selection $R^{\prime}$ of size at most $m(k,d)$ from the families $\mathcal{F}_{1}^{\prime},\dots,\mathcal{F}_{d+(d-k)+1}^{\prime}$ contains a $k$-dimensional cone with an apex at the origin. On the other hand, for every $i\in[d+(d-k)+1]$, the intersection set of $\mathcal{F}_{i}^{\prime}$ can be described as the solution set of a finite homogeneous system of linear inequalities $\left\langle a_{i},x\right\rangle\leq 0$, $a_{i}\in\mathcal{A}_{i}\subset{\mathbb{R}}^{d}\setminus\{0\}$. The sets $\mathcal{A}_{1},\dots,\mathcal{A}_{m(k,d)}$ satisfy the assumption of Theorem 1.3. Thus, for some $i\in[d+(d-k)+1]$, the system $\left\langle a,x\right\rangle\leq 0,\;a\in\mathcal{A}_{i}$, has $k$ linearly independent solutions, that is, the intersection $\bigcap\limits_{F_{i}^{\prime}\in\mathcal{F}_{i}^{\prime}}F_{i}^{\prime}$ contains a $k$-dimensional cone with an apex at the origin. As shown, this means that the intersection $\bigcap\limits_{F_{i}\in\mathcal{F}_{i}}F_{i}^{\circ\circ}$ contains a $k$-dimensional cone with an apex at the origin.
Applying Lemma 6.5, we see that each of the sets of $\mathcal{F}_{i}$ contains a translate of some $k$-dimensional cone. The proof of Theorem 6.3 is complete. ∎
We believe that, under the assumptions of Theorem 6.3, there is a family whose intersection set contains a $k$-dimensional cone.